(1) A sphere of decaying radioactive material of radius ro produces heat at a rate of q (W/m3). The sphere is contained in
Answer #1

a) for the radioactive material, consider a spherical shell of thickness dr at a radial distance of r (0<r<ro) from the center

Applying heat balance on this particular shell:

In- Out GenerationAccumulation

Assuming steady state, Accumulation =0

OT Or

Out =-K_{o}A\frac{\partial T}{\partial r}|_{r+dr}=-K_{o}(4\pi r^{2})\frac{\partial T}{\partial r}|_{r+dr}

, Generation-q(4π , 2.Sr )

от ar Эт

Dividing by \Delta r and \Delta r\rightarrow 0

or

where q''' is the heat generated per unit volume of radioactive material

For the graphite shell, the equation remains same except the generation term changes. Since there is no generation inside this shell,

Generation-0

Hence the energy balance for ro<r<r1,

ar ar

Dividing by \Delta r and \Delta r\rightarrow 0

1HTr Эт

b) The boundary conditions,

By Energy transferred at the interface should be balanced, at r=ro

1.от o ar at r=ro

2. At r=r1, T=T\infty at steady state